### v1

# Object as a Service: Simplifying Cloud-Native Development through Serverless Object Abstraction

Abstract
The function-as-a-service (FaaS) paradigm is envisioned as the next generation of cloud computing systems that mitigate the burden for cloud-native application developers by abstracting them from cloud resource management. However, it does not deal with the application data aspects. As such, developers have to intervene and undergo the burden of managing the application data, often via separate cloud storage services. To further streamline cloud-native application development, in this work, we propose a new paradigm, known as Object as a Service (OaaS) that encapsulates application data and functions into the cloud object abstraction. OaaS relieves developers from resource and data management burden while offering built-in optimization features. Inspired by OOP, OaaS incorporates access modifiers and inheritance into the serverless paradigm that: (a) prevents developers from compromising the system via accidentally accessing underlying data; and (b) enables software reuse in cloud-native application development. Furthermore, OaaS natively supports dataflow semantics. It enables developers to define function workflows while transparently handling data navigation, synchronization, and parallelism issues. To establish the OaaS paradigm, we develop a platform named Oparaca that offers state abstraction for structured and unstructured data with consistency and fault-tolerant guarantees. We evaluated Oparaca under real-world settings against state-of-the-art platforms with respect to the imposed overhead, scalability, and ease of use. The results demonstrate that the object abstraction provided by OaaS can streamline flexible and scalable cloud-native application development with an insignificant overhead on the underlying serverless system.
Index Terms:
FaaS, Serverless paradigm, Cloud computing, Cloud-native programming, Abstraction.
I Introduction
I-A FaaS and Its Shortcomings
Function-as-a-Service (FaaS) paradigm is becoming widespread and envisioned as the next generation of cloud computing systems (a.k.a. Cloud 2.0) [18] that mitigates the burden for programmers and cloud solution architects. Major public cloud providers offer FaaS services (e.g., AWS Lambda, Google Cloud Function, Azure Function), and several open-source platforms for on-premise FaaS deployments are emerging (e.g., OpenFaaS, Knative). FaaS offers the function abstraction that allows users to develop their business logic and invoke it via a predefined trigger. In the backend, the serverless platform hides the complexity of resource management and deploys the function seamlessly in a scalable manner. FaaS is proven to reduce development and operation costs via implementing scale-to-zero and charging the user in a truly pay-as-you-go manner. Thus, it aligns with modern software development paradigms, such as CI/CD and DevOps [7].
As the FaaS paradigm is primarily centered around the notion of stateless functions, it naturally does not deal with the data. However, in practice, most use cases must maintain some form of (structured or unstructured) state and keep them in the external data store. Thus, often the developers have to intervene and undergo the burden of managing the application data using separate cloud services (e.g., AWS S3 [4]). Even though stateless functions make the system scalable and manageable, the state still exists in the external data store, and the developer must intervene to connect the function to the data store. For instance, in a video streaming application [12], developers must maintain video files, metadata, and access control in addition to developing functions.
Apart from the lack of data management, current FaaS systems do not offer any built-in semantics to limit access to the functions’ internal (private) mechanics. Providing unrestricted access to the developer team has known side effects, such as function invocation in an unintended context and data corruption via direct data manipulation. To overcome such side-effects, developers again need to intervene and undergo the burden of configuring external services (e.g., AWS IAM [2] and API gateway [1]) to enable access control.
Last but not least, current FaaS abstractions do not natively support function workflows. To form a workflow, the developer has to generate an event that triggers another function in each function. However, configuring and managing the chain of events for large workflows becomes cumbersome. Although function orchestrator services (e.g., AWS Step Function [3] and Azure Durable Function [9]) can be employed to mitigate this burden, the lack of built-in workflow semantics (see Figure 1) in FaaS forces the developer to intervene and employ other cloud services to navigate the data throughout the workflow manually. In sum, although FaaS makes the resource management details (e.g., auto-scaling) transparent from the developer’s perspective, it does not do so for the data, access control, and workflow management.
I-B Proposed Paradigm
(a) Function as a Service (FaaS)
(b) Object as a Service (OaaS)
Figure 1: A bird-eye view of FaaS vs. OaaS.
To overcome these inherent problems of FaaS, we propose a new paradigm on top of the function abstraction that mitigates the burden of resource, data, and workflow management from the developer’s perspective. We borrow the notion of “object” from object-oriented programming (OOP) and develop a new abstraction level within the serverless cloud, called Object as a Service (OaaS) paradigm.
As shown in Figure 1, unlike FaaS, OaaS segregates the state management from the developer’s source code and incorporates it into the serverless platform to make it transparent from the developer’s perspective. OaaS also incorporates workflow orchestration via its dataflow abstraction with built-in data navigation across functions. Incorporating the application data and workflow into the object abstraction unlocks opportunities for built-in optimization features, such as data locality, data reliability, caching, software reusability [13], and data access control. OaaS provides an interface for developers to declare the behavior and properties of objects in the form of class and function. For more complex use cases where the developer needs to orchestrate multiple objects and functions in the form of a workflow, OaaS offers the native (built-in) dataflow semantic [37] that can hide details of the data navigation and synchronization via declaring the flow of data in the form of an object’s function.
An exemplar use case that can take advantage of OaaS is a cloud-based video streaming system (e.g., [12, 31]) that needs developers to implement new streaming services for the available video content rapidly. Examples of such services are blurring harmful (gore) content from videos, detecting faces on surveillance videos, and making the video processing pipeline for transcoding. Implementing these services using FaaS entails dealing with the state data (i.e., videos), in addition to developing the business logic. In this scenario, the OaaS paradigm can mitigate the developer’s burden by offering encapsulation. The videos are defined as persistent objects bound to a set of functions that can be invoked by the viewer’s application and potentially change the object’s (video’s) state. For instance, to blur gore segments in the video object v1, the developer invokes v1.detect_gore() to get output object g1 that contains a list of time stamps of the gore segments. Next, blurring the video is performed via invoking v1.blurs(g1).
I-C Research and Contributions
The first and foremost challenge to establish OaaS is to offer object abstraction with an interface for developers to declare the behavior and properties of objects in the form of class, function, and dataflow semantics. Together with access modifiers, OaaS can hide the internal mechanics (i.e., private function) from other developers and users. The object can have pointers to other objects and form a dataflow function on top of them. Such a high-level object can hide the details of low-level ones by declaring access modifiers that reject any function calls that involve invalid access.
To realize the OaaS paradigm, we develop the Oparaca (Object Paradigm on Serverless Cloud Abstraction) platform with the concern of the flexibility and modularity, so that it can accommodate different use cases. The second challenge is to design Oparaca to be pluggable and can work with different execution (i.e., FaaS) and storage modules. We avoid having Oparaca tightly depend on the execution module via exploiting the notion of pure function that offloads the task to the execution module without any side-effect. We also devise a data tiering scheme as the abstraction layer between the storage and execution modules. Therefore, changing the storage type does not imply changing the function code.
Although utilizing the aforementioned invocation scheme provides Oparaca with ease of use and flexibility, it causes additional overheads, mainly because of moving the data between OaaS modules. This is particularly important for unstructured data, also known as BLOB (e.g., multimedia content), that usually persists on a separate storage type, such as object storage. Therefore, the third challenge is to design Oparaca so that supporting objects’ overhead is minimal and tractable. To mitigate the overhead, the data tiering scheme within the Oparaca platform diminishes the latency of accessing the object by using distributed in-memory hash table [40]. To further mitigate the overhead of accessing objects, Oparaca is equipped with a presigned URL with a redirection mechanism that reduces unnecessary data movements within the platform instead of relaying (transferring) the object state.
In OaaS, concurrent access requests can be made to the same state, which makes it prone to data inconsistency from race conditions. Moreover, any system failure can potentially cause state inconsistency between multiple data stores. For example, if the function updates the state to the first storage but fails to update the second storage, it may cause an inconsistency. Thus, the fourth challenge is to make Oparaca fault-tolerant and maintain data consistency across multiple data stores. To overcome this challenge, we develop a fail-safe state transition in Oparaca to maintain consistency across various data stores. In addition, we develop a method based on the lightweight optimistic locking [29] together with localized locking to avoid concurrent same-object modification. Another aspect of fault tolerance is the recovery mechanism from failure via retrying. However, such a remedy by itself can lead to another problem—repeating the execution more than once and falling into an undesirable state. To cope with this challenge, we develop a mechanism within Oparaca that establishes consistent state transitioning by guaranteeing “exactly-once” execution for the function calls.
In summary, the key contributions of this research are as follows:
  1. 1.
    Proposing the OaaS paradigm to hide data, workflow, and resource management complexities from the user view, and to mitigate the redundancy complexity through developing the cloud object inheritance mechanism.
  2. 2.
    Developing a prototype of OaaS, called Oparaca platform [^1] The source code, packages, and example use cases are available here: https://github.com/hpcclab/OaaS, that can support both structured and unstructured types of states while maintaining data consistency and fault-tolerance with minimal performance overhead.
  3. 3.
    Minimizing the overhead of Oparaca via presigned URL with redirection and caching mechanisms.
  4. 4.
    Analyzing Oparaca’s performance from the scalability, overhead, and ease-of-use perspectives.
In the rest of this paper, Section II reviews the state of the art in the serverless paradigm. Section III discusses the conceptual design. Section IV presents the architecture of Oparaca and internal details. Section V discusses non-focused topics in this work. Section VI evaluates the overhead, scalability, and development efficiency of OaaS. Finally, we conclude this paper in Section VII.
II Background and Prior Studies
II-A FaaS Paradigm
The FaaS paradigm allows the developer to implement the application as a set of independent functions that are transparently provisioned in isolation on the cloud infrastructure. FaaS is offered by public cloud providers (e.g., AWS Lambda, Azure Function, Google Cloud Function).
FaaS can also be self-hosted via open-source platforms (e.g., OpenFaaS [14] and OpenWhisk [15]). FaaS invokes the function that matches its predefined trigger(s) upon receiving the event.
A variant of FaaS, Container as a Service (CaaS) [20], does not offer the function development framework. Instead, the user must provide the already-containerized function. Kubernetes [10] is a widely-used platform that automates container provisioning (via YAML configuration language) and manages the life-cycle of containerized services. Knative [16] complements Kubernetes by enabling FaaS features at the container level. It is composed of two main components: Knative Serving, and Knative Eventing. The former enables auto-scaling, scale to zero, and minimal configuration of the containerized services. The latter enables pipelining and routing events to streamline the development of event-driven containers.
II-B State Management Strategies
The idea of stateful serverless is explored in several research works. As noted in Figure 2 and Table I, these works can be categorized into actor model, datastore abstraction, and pure function approaches depending on where the platform stores the state data, and how the function accesses the data. According to Figure 2, the actor model places the state inside the worker node (instance) to achieve the data locality. In the pure function, however, the state is placed on other systems (e.g., database) and is transferred to the worker instance upon invocation. Hence, the state appears as part of the function input argument, and the modified state appears as its output. Thus, the function is still stateless while exhibiting stateful features. The datastore abstraction is a hybrid approach where the platform provides the API for the function to access the data on demand. Depending on the design, the state might be stored in the database but can also be cached in the worker.
Figure 2: The illustrated comparison of three different models of stateful serverless.
Actor
Model
Datastore
Abstraction
Pure
Function
Data
placement
worker
instances
platform services
or database
(depend on impl.)
platform
services
or database
Complexity high depend on impl. low
Data locality high depend on impl. low
Unstructured
data support
difficult yes difficult
Deployment
granularity
actor
(multiple
functions)
function function
Maintain-
ability
low high high
Solutions
Kalix [22],
Azure Entity
Func.[33],
Cepless [32]
Cloudburst [38],
FAASM [35],
Apiary [27],
Crucial [8],
Boki [25],
Beldi [44],
Oparaca
Kalix [22]
Statefun [5],
Oparaca
TABLE I: Comparing properties of various design patterns to build a stateful serverless platform.
According to Table I, the actor model serverless platform needs to maintain the availability of each actor (object) where both data and compute reside. Maintainability is particularly difficult for bulky unstructured data because the platform needs to balance each node’s computing and storage aspects. In addition, the platform has to support a routing mechanism to navigate a function call to the actor’s location. Alternatively, the pure function approach disaggregates the state management and compute (function) for the sake of system design simplicity. However, it compromises the data locality aspect. Similar to pure functions, datastore abstraction also relaxes the need to store the state on the worker node. Regardless, it utilizes caching techniques to preserve data locality. The deployment granularity of the actor model approach is an actor with multiple functions that share the same state, whereas the granularity in other approaches is a single function.
The actor model approach has been popular in programming languages and OOP, because it spurs asynchronous messaging across actors and lends itself to distributed deployments. That is why it has been an attractive choice for stateful serverless platforms, even though it poorly supports unstructured data. Kalix [22] and Azure Entity Functions [33], which are part of Azure durable functions, are example platforms implemented based on the actor model approach. The serverless platforms based on the datastore abstraction are mostly popular in research. Cloudburst [38] offers stateful functions using a shared distributed key-value database to keep track of the state. FAASM [35] optimizes the function-state interaction overhead via employing web assembly [17]. Even though web assembly enables multiple functions to share the memory and achieve data locality, it implies compiling the code into web assembly, which limits the compatibility of some libraries. Crucial [8] allows a function to access the shared data via the DSO layer (distributed hash table). Boki [25] enables stateful function via providing API access to the distributed logging system. Beldi [44], on the other hand, provides the database and transaction API to the state. Apache Flink Stateful Function (StateFun) [5], and Kalix [22] are solutions based on the pure function approach. Statefun is built atop Apache Flink based on the actor model. However, it offloads the function code to a dedicated worker that can be categorized under the pure function approach.
III Object as a Service (OaaS) Paradigm
III-A Conceptual Modeling of OaaS
To realize OaaS, first, we need to establish the notion of cloud object as an entity that possesses a state (i.e., data) and is associated with one or more functions. We empower objects to support both structured (e.g., JSON records) and unstructured (e.g., video) forms of state. Upon calling an object’s function, OaaS creates a task that can safely take action on the state.
Second, OaaS provides the class semantic as a framework to develop objects. Inspired by OOP, the developer has to define a set of functions and states within the class. Then, an arbitrary number of objects—that is bound to the functions and states declared in that class—can be instantiated. To improve cloud software reusability and maintainability, we enable class inheritance for cloud functions and states from other classes, plus the ability to override any derived function.
Third, OaaS offers built-in access control to provide the ability to declare the “scope of accessibility” for a state or function. Importantly, to define a set of related classes, the developer can declare it within a single module (a.k.a. package) that includes the access modifier to prevent unauthorized access to the state(s)/function(s) from other packages. This is particularly useful when cloud application developers utilize imported third-party packages.
Fourth, to enable higher-level abstractions for the users or developers, OaaS allows combining (nesting) cloud objects into one. The high-level object holds a reference to the lower-level object(s), and the invoked function can leverage the reference to fetch the lower-level object as its input. Importantly, we make it possible that the high-level object implements a dataflow function (called macro function); i.e., , invoking a chain of functions from the lower-level objects. This resembles configuring a workflow in conventional FaaS [3], albeit in a transparent manner. The major difference between macro functions and other FaaS workflows is that macros introduce the flow of execution via the flow of data rather than the invocation order. That is, given the dataflow semantics, the execution flow in a macro function is determined by the flow of data, and the developer only needs to introduce the flow. Then, in the background, OaaS handles the parallelism and data navigation while guaranteeing state consistency.
Listing 1: An example simplified script that declares multimedia package with a video class, and a transcode function for it in the YAML format.
1name multimedia
2classes
3 - name video
4 stateSpec
5 keySpecs
6 - name mp4
7 access PUBLIC
8 functions
9 - function new
10 access PUBLIC
11 - function transcode
12 access PUBLIC
13 outputCls .video
14functions
15 - name transcode
16 type TASK
17 image transcode-pylatest
18 ...
III-B Developing Classes in OaaS
In OaaS, developers define one or more classes within a package using configuration languages like YAML or JSON. The package definition contains the class section and the function section. The functions section defines the configuration and deployment details of each function. The class section defines the object’s structure, which includes the state and function it links to.
Figure 3: Different types of functions supported by OaaS.
As shown in Figure 3, OaaS supports three function types. First, built-in functions that are provided by the platform. These functions could be the standard functions such as CRUD (create, read, update, and delete), which are the common data manipulation operations. The platform manages the execution of these functions without intervention from the developer. Second, custom (a.k.a. task) functions that are developed by developers (OaaS users) to provide their business logic. To handle the invocation of these functions, OaaS employs existing FaaS engines in its underlying layers to exploit their auto-scaling and scale-to-zero capabilities. Third, dataflow (macro) functions are defined as a DAG representing the chain of invocations to objects.
As an example of package definition, Listing 1 represents a declaration example for a package that includes one class called video that has a state named mp4 (Line 6), built-in function named new (Line 9), and custom function named transcode (Line 1). The state mp4 refers to video data that is unstructured data. The class has a public custom function called transcode. The definitions of the custom function are declared in Lines 15—17. The type of a function (Line 16) can be a task (or a macro, as noted earlier). This function creates another object instance of type video as an output. Line 17 declares the container image URI for executing function code.
IV Oparaca: A Platform for the OaaS Paradigm
IV-A Design Goals
Now that the notion of OaaS is established, we elaborate on the design goals and architecture of the Oparaca platform. We set two design goals for Oparaca:
The first goal considers the fact that even though OaaS provides a higher-level abstraction for cloud-native application development, it is not always a replacement for FaaS. Hence, Oparaca should be backward compatible so that the developer can still use stateless FaaS and retrieve the data from the storage system directly. Additionally, most FaaS solutions support both synchronous and asynchronous function invocations. Therefore, Oparaca should support both types of function invocations as well for backward compatibility.
Second, for the sake of extensibility, Oparaca separates the components of the control plane from the execution plane, and the execution plane can be performed without the knowledge of its control plane but communicate through the standardized APIs. In this manner, the design is platform-agnostic and is not bound to a single preset underlying FaaS system, as its execution plane. This enables Oparaca to accommodate various types of execution planes, optimized for the requirements of different use cases, such as supporting latency-constrained function calls [36]; or supporting access to hardware accelerators [41].
IV-B Overview of the Oparaca Architecture
Figure 4: A bird-eye view to the architecture of Oparaca. Dashed lines show the workflow of actions initiated by the developer to define classes and objects, and solid lines show them for the end-user to make use of objects and invoke their functions either synchronously (red arrows) or asynchronously (blue ones).
The Oparaca platform is designed based on multiple self-contained microservices that communicate within a serverless system. Figure 4 provides a birds-eye view of the Oparaca architecture that is composed of five modules:
  • Class Module serves as the interface for developers to create and manage classes and their functions.
  • Object Module serves as the cornerstone of Oparaca that has two main objectives: (a) providing the “object access interface” for the user application to access an object(s); and (b) offering the object abstraction while transparently handle function invocation and state manipulation.
  • FaaS Engine is the underlying execution plane of Oparaca, which can be any of the existing FaaS systems (e.g., Knative).
  • Data Management Module that handles the objects’ state.
  • Ingress Module whose purpose is to provide a single end-point for the user application.
Details of these modules, their interactions, and how they fulfill the consistency and fault-tolerance objectives (described in Section  I) are elaborated in the following subsections.
IV-C Class Module
To define a new class and its functions in Oparaca, the developer must bundle them as one package and register them to the Package Manager, shown in Figure 4. Upon successful package validation by the Package Manager, the Class Runtime Manager (termed CRM for brevity) performs the class registration process that includes two operations:
(a) Informing the Object Module about the new/updated class. Upon receiving a class registration request, the Object Module creates a handler instance to be prepared for handling object invocation requests. We elaborate on this process in Section IV-D.
(b) Registering the custom functions of the new class in the FaaS engine, so that they can be invoked at a later time. Recall that we aim to make Oparaca agnostic from the underlying FaaS engine. Therefore, the function registration process should be doable for various FaaS engines. To that end, we design Oparaca to be able to host a dedicated CRM for each FaaS engine. Therefore, a new FaaS engine can be integrated into Oparaca by simply plugging its dedicated CRM into the system. Upon receiving the function registration event, only the specific CRM associated with the configured FaaS engine processes the event, which includes translating the developer-defined function configuration into the FaaS-engine-specific format and then forwarding it to the associated FaaS engine (e.g., Knative). Consequently, the underlying FaaS Engine creates the actual function runtime to be invoked by the Object Module.
IV-D Object Module and FaaS Engine
Figure 5: The cluster of Invokers replicates and distributes object data across the cluster memory via consistent hashing technique and in-memory data grid (IMDG) to reduce the data transfer overhead. The invoker offloads the invocation to a corresponding FaaS engine function that must pass through the reverse proxy.
Recall that OaaS needs to support three types of functions: built-in, custom, and dataflow. Unlike built-in functions (those provided by the platform) and dataflow functions (those to handle developer-defined DAGs) that can be executed without the direct need of the FaaS engine, custom functions need to execute the developer-provided code on the FaaS engine. Thus, Oparaca requires a mechanism to utilize the FaaS engine to execute the custom function code while allowing it to access the object state transparently and with the minimum data transfer overhead. Needless to say, this mechanism also maintains the separation between the Object Module and the FaaS engine.
To fulfill the aforementioned expectations, in the Object Module, we design the object invocation mechanism to be cognizant of the type of state (structured vs unstructured) and manage it such that the data transfer overhead needed for the invocation is minimized. We develop a hybrid approach that leverages the “pure function” technique to handle invocations with structured data and the “datastore abstraction” technique to handle cases with unstructured data. The rationale of this design choice is that the unstructured state (i.e., BLOB) is usually large and expensive to transfer, hence, to maintain efficiency, the FaaS engine should retrieve the state directly from the object storage (e.g., S3) in a lazy (on-demand) manner. This differs from the structured state for which we include the state as an input argument to maintain a clear separation between the Object Module and the FaaS engine and let the FaaS engine maintain its statelessness.
In the Oparaca architecture (Figure 4), the described invocation and state handling mechanism is handled by the Invoker component. In particular, to offload the object invocation to the FaaS engine, Invoker bundles the request and the related structured object data as a “task”, as described in the next part, and passes it to the associated FaaS engine for execution.
IV-D1 Task Generation in the Invoker Component
To further reduce the data transfer overhead of providing the object abstraction in the task generation process, we design Invokers to maintain the object data (i.e., state and metadata) in a distributed hash table [19], thereby reducing the cost of data transfer in a scalable manner. As shown in Figure 5, we equip each Invoker instance with an embedded in-memory data grid (IMDG) [43]. IMDG partitions the entire DHT into multiple segments and distributes them across Invoker instances. The invoker with IMDG determines the segment for a given object by consistent hashing of the object ID and assigns the object data to the selected segment. Similarly, to retrieve the object data, IMDG determines the object data location and then fetches it from the owner of the segment in at most one hop. Upon receiving a function call, the Invoker bundles the invocation request and associated object data into the task and offloads it to be executed on the FaaS engine.
IV-D2 Unstructured Data Accessing
To minimize the overhead of accessing unstructured data, Oparaca allows function code to access the unstructured data on-demand and directly through a presigned URL and redirection mechanism. The presigned URL is the specific HTTP URL that includes the digital signature in query parameters to grant permission for anyone with this URL to access the specific data without the secret token. When a function needs to access the unstructured data, it sends an HTTP request to the storage adapter to receive the redirection response that points to the presigned URL of specific state data. Then, the function code can fetch the content directly from object storage via the given presigned URL. In addition to minimizing the overhead, using the presigned URL is important in protecting the function container from unauthorized access to other objects’ data by analyzing their URL patterns.
IV-D3 Task Completion
After the FaaS engine completes the task, it sends the task completion data to the Invoker to update the state. If the function reports a failed task, the state remains unchanged. Otherwise, the Invoker updates the object data in IMDG and then writes it to the persistent database immediately or asynchronously. If an invocation involves both structured and unstructured states, we use pure and datastore techniques together, which can potentially lead to ”state inconsistency challenges”. We address this challenge in section IV-F.
IV-D4 Synchronous and Asynchronous Invocation
As mentioned in Section IV-A and shown in Figure 5, we designed Oparaca to offer two modes of function invocation: synchronous and asynchronous. In the synchronous mode, the function is executed immediately upon invocation and returns the result to the caller. Meanwhile, in the asynchronous mode, the invocation ID is provided to the caller as a reference, so they can later check the invocation result. The request is placed into the message broker to be reliably processed at a later time. To accommodate both modes, the Invoker utilizes the handler instance to accept the invocation request for either the REST API (synchronous) or the message broker (asynchronous). Subsequently, the handler instance forwards the request to be processed in the same way by the other part of the Invoker.
Figure 6: The process of offloading invocation task into the function runtime. Invoker bundles the request input and object state into a task and offloads it to the function to be executed. With fail-safe state transition, when the function needs to update the file in object storage, it creates a new file and updates the corresponding version ID via the returning completion message.
IV-E Ingress Module
To provide the end user with a single access point, we position the Ingress Module in front of the cluster of Invokers. Additionally, to minimize data movement, the Ingress Module is designed to be aware of the object data distribution through consistent hashing of DHT. This allows the Ingress Module to correctly forward the object invocation request to the Invoker that owns the primary object data. As a result, the designated Invoker is able to access the data in its memory.
IV-F Resilience Measures of Oparaca
Oparaca is prone to the data inconsistency problem that stems from both failure and race conditions. In this section, we describe the internal mechanisms of Oparaca designed to make it resilient against these conditions.
IV-F1 Resilience against failure
Data inconsistency from failure can happen if the system stops while performing multiple update operations, causing some of the update operations to be incompletely executed. The pure function model, used for structured data in Oparaca, is inherently immune to this problem because a function returns the modified state to the platform only when its execution is complete. Nonetheless, the datastore abstraction used for the unstructured data in Oparaca is still prone to the data inconsistency problem between the structured database and object storage.
Maintaining data consistency across two data storages implies guaranteeing both storages are either successfully updated or fail for the same invocation. Otherwise (i.e., if only one of them succeeds), it leads to data inconsistency. To overcome this problem, we develop the fail-safe state transition mechanism that disregards the data update in the object storage if Invoker fails to update the structured part of the object data in the structured database. For that purpose, the mechanism uses a two-phase versioning scheme to keep track of the unstructured data. As shown in Figure 6, in the first phase, the mechanism creates a version ID for each file (unstructured data) and keeps them as structured data (metadata of object data) to track the current version of the file. In the second phase, which occurs upon function completion, Invoker changes all version IDs associated with the updated files (unstructured data) and then writes them to IMDG and the structured database.
For example, consider object $o_{1}$ that has file $f_{1}$ with the version ID $v_{1}$. Upon function invocation, $f_{1}$ is updated and written to the object storage with version ID $v_{2}$. After the execution, the Invoker must change the version ID from $v_{1}$ to $v_{2}$ and commit the new structured object data. If any operation fails within this process, the next invocation still loads $o_{1}$ with version ID $v_{1}$, as if the previous invocation never happened. In the last step, when the invocation is complete, the Invoker purges the old and unused versions of data.
IV-F2 Resilience against race condition
Race conditions in Oparaca can occur when multiple invocations modify the same object data simultaneously, resulting in potential data inconsistency. One way to prevent this issue is by using database transactions; however, this method lacks abstraction as it allows direct function code access to the database and is tightly dependent on the type of database. An alternative approach to avoiding race conditions is the cluster-wide pessimistic locking mechanism to synchronize the locking state for all invokers. Nevertheless, this approach necessitates additional network communication to coordinate the locking state, which can lead to scalability issues. Alternatively, we develop an improved version of this mechanism, called ”localized locking,” which relies on consistent hashing to direct the invocation request to the invoker that owns the primary copy of the targeted object data. Each invoker will only need to lock the object locally without additional network communication, making it more scalable than the cluster-wide version. Additionally, our localized locking approach guarantees that requests to the same object are executed in the arrival order, which is necessary in certain use cases where order matters, such as seat reservations. This is difficult to achieve with cluster-wide locking.
IV-F3 Failure recovery in Oparaca
To further establish resilience against failures, Oparaca is equipped with a mechanism to self-recover from the failure. Broadly speaking, a function invocation failure can be recovered by simply retrying the invocation. However, this approach can cause data incorrectness owing to the execution of the function more than once. The retrying approach could be undesirable for synchronous invocations because the failure can be handled on the client side. For asynchronous invocations, however, we need to guarantee that any invocation is only executed exactly once.
To achieve the exactly-once guarantee, we have to prevent three sources of the problem that are: (a) losing messages, (b) duplicating messages, and (c) processing messages more than once. Message brokers with stable storage (e.g., Kafka [28]) have features that can be leveraged to address these problems. To solve the first problem, upon failure occurrence, the Invoker can detect and reprocess the incomplete request using an offset number that is automatically generated by the message broker. The offset number is the auto-incremental number based on the message’s arrival order and can be used to track the message’s position in the queue. The second problem of producing duplicated request messages can be resolved using the message broker’s “idempotent producer” feature.
However, the message broker cannot completely address the third problem. That is, the Invoker can process the same invocation request more than once when the message broker has not acknowledged the completed one before the system failure occurs. We prevent this problem by tracking the offset number of the last processed request and adding it to each object metadata. In this manner, before processing an invocation request, Invoker checks the offset number of the target object to see if it is lower than the offset number of an incoming request. When the condition is met, the Invoker can detect it has not been processed and perform the normal operation. Otherwise, it must be skipped to avoid reprocessing.
IV-G Dataflow Abstraction in Oparaca
To offer a higher level abstraction to declare a workflow of invocations, Oparaca provides the dataflow (as opposed to the standard workflow) abstraction as a built-in feature to enable developers to define the invocation steps in the form of a directed acyclic graph (DAG). In every step, the developer can declare the output of each invocation as a temporary variable within the workflow. Then, the next invocation can use the temporary variables from previous steps as the input or target to call the function. Upon the developer registering a dataflow function, Invoker constructs the DAG by having the invocation step as the edge and the objects as nodes.
When a request for dataflow execution is received, one of the Invokers takes on the role of orchestrator, similar to the orchestrator pattern [34] in microservices. It breaks down the dataflow into multiple lower-level invocations and forwards them based on the topological order of DAG. Using consistent hashing, the invoker can determine the address of the target object and send the request directly to another Invoker that holds the target object. When each step is completed, the orchestrator keeps track of the intermediate dataflow state to transparently operate the data exchange between invocation steps. With the orchestrator pattern, the dataflow control logic is centralized into a single invoker, simplifying the management, monitoring, and error-handling implementation.
When using the orchestrator pattern, the exact-once guarantee may be compromised because the object data is stored separately from the dataflow state. If the guarantee is needed, Oparaca allows developers to mitigate this problem by flagging all invocation steps on as immutable. Upon handling the dataflow request, Oparaca can generate the output ID in advance for each step, making each step of dataflow execution idempotent and safely re-executable.
V Discussion
V-1 Security
Certain security measures are put in place in Oparaca to strengthen it against potential attacks. The first measure is to reduce the attacking surface by limiting the necessary outbound traffic from the function runtime. As the function container only requires access to the storage adapter and object storage, the traffic policy can be configured to block outbound traffic except for the Storage Adapter and object storage. The second measure is to avoid reusing secret tokens. To prevent the function container from accessing out-of-context data via analyzing the URL path, we use the presigned URL mechanism for object storage. Thus, object storage in Oparaca is more secure than in FaaS, where the same secret key is used for every request. To secure the storage adapter, we can make the Invoker generate a unique secret token for each task, and every request for the storage adapter must be authenticated via the secret token.
V-2 Cold Start
The developer functions and the Oparaca components can benefit from scale-to-zero to reduce the cost when there is no usage. However, this has the side-effect of more cold starts. Since Oparaca components are shared across functions, we can effectively keep it warm to eliminate the additional cold start impact. In such a case, the cold start performance entirely depends on the underlying serverless execution engine.
VI Performance Evaluation
(a) Video transcoding (sync)
(b) Text concatenation (sync)
(c) JSON update (sync)
(d) Video transcoding (async)
(e) Text concatenation (async)
(f) JSON update (async)
Figure 7: The average execution time of functions for objects with various state sizes in synchronous and asynchronous invocations. Two versions of Oparaca are examined: the full version and without URL-redirection version (oprc-relay). We also capture the time used by the internal Knative part in both Oparaca versions and show them with suffix -exec and plot it in the same bar as its Oparaca version.
VI-A Experimental Setup
We deploy the Oparaca platform on 4 machines of Chameleon Cloud [26], each with 2 sockets of 24-Core Intel(R) Xeon(R) Gold 6240R CPU processors that collectively have 192 cores, 768 GB memory, and SSD SATA storage. We set up the Kubernetes cluster, which includes 15 VMs with 16 vCPUs and 32 GB memory. We made another 2 VMs for the S3-compatible storage (Minio [23]) for unstructured data and ArangoDB ([21]) for structured data. Oparaca is implemented using Java [6] with Infinispan [24] for IMDG.
Baselines. We configure Apache Flink Stateful Function (StateFun) [5], OpenWhisk [15], and Knative [16] to serve as the baselines. Unlike Oparaca and OpenWhisk, which focus on API calls and event handling, StateFun is an open-source stateful serverless system focusing on stream processing. Because StateFun does not manage the function worker instances out of the box, we configure Knative to complement it. OpenWhisk and Knative are popular open-source stateless FaaS platforms that we use to represent the state management done by the developer.
We used Gatling [11] for load generation and implemented three applications to serve as the workload. First is the video transcoding function using FFmpeg [42], which is a CPU-intensive application. The second is a text concatenation function that concatenates the content of a text file (state) with an input string. This function represents a highly IO-incentive workload. Third is the JSON update function, which has only structured data as a JSON and is used for randomly putting key-value pairs into the JSON state data. The rest of the workload characteristics are specific to each experiment and are explained in the respective sections. All three functions are implemented in Python language [39].
VI-B Analyzing the Imposed Overhead of Oparaca
The abstractions provided by Oparaca are not free of charge and introduce some time overhead to the applications using these abstractions. In this experiment, our aim is to measure this overhead and see how the efficient design of Oparaca can mitigate this overhead. The latency of a function call is the metric that represents the overhead. We mainly study two sources of the overhead: (a) The state data size that highlights the overhead of OaaS in dealing with the data, and (b) The concurrency of function calls that highlights the overhead of the Oparaca system itself.
(a) Video transcoding function (sync)
(b) Text concatenation function (sync)
(c) JSON update function (sync)
(d) Video transcoding function (async)
(e) Text concatenation function (async)
(f) JSON update function (async)
Figure 8: The average completion time of functions upon varying the rate of incoming requests in synchronous and asynchronous invocations. oprc-queue is the queuing time that requests stay within the message queue
The impact of changing the state size is shown in Figure 7. To generate objects with various state sizes, we increased the input video length from 1—30 seconds. To remove the impact of video content on the result, the longer videos were generated by concatenating the same 1-second video. Similarly, the text files are from 0.01—16 MB. For the JSON object, the key and value sizes are 10 and 40 bytes, respectively, and the number of key-value pairs varies from 10—320 pairs. To concentrate only on the overhead of data access and avoid other sources of overheads, we configure Gatling to assign only one task at a time and set it to repeat this operation 100 times. To analyze the improvements offered by the URL redirection, we examine two versions of Oparaca: the full version (expressed as oprc) and without URL redirection (expressed as oprc-relay). The horizontal axes represent different state sizes for video, text, and JSON, respectively, and the vertical axes represent the average response/completion time (latency).
In Figure 7, the average task execution time increases for larger state sizes. For the video transcoding function, all of the platforms perform with similar latency, which is expected because of the compute-intensive nature of the video transcoding that dominates the completion time. In a text concatenation function, however, Knative performs slightly better than Oparaca because of the overhead of unstructured state access by the redirection of the presigned URL. However, if we compare Oparaca with another version that uses a relay mechanism to provide the state abstraction, it performs much lower than its alternative with an average of 30% lower response time. Lastly, we can see all the described trends happen similarly for synchronous and asynchronous request types.
In the JSON update function (Figures 7c and 7f), Oparaca can perform with lower latency than Knative because the function does not need to fetch the object data from the database because of the pure function semantic. Nevertheless, Knative can catch Oparaca by increasing the key-value entries to 320. The reason is that the gain from eliminating the database connection is surpassed by the overhead of moving the data to the function code for larger records. OpenWhisk and Knative have the same pattern because both of them are FaaS, but OpenWhisk performs significantly worse. In Figure 7f, the Statefun shares the same pattern with Oparaca with the consistent gap because it also relies on local storage to keep the function state without the need to fetch the data from the database. We also observe that Statefun performance degraded compared to our initial results [30]. This is because the storage hardware being used for the experiment has a lower through, which impacts the performance of Statefun.
(a) Speedup results by horizontal scaling
(b) Throughput results from horizontal scaling
Figure 9: Evaluating the scalability of the OaaS platform against other baselines.
Figure 10: Evaluating the performance of localized locking compared to cluster-wide locking
The impact of concurrent function invocations on the Oparaca overhead is shown in Figures 8. In synchronous invocation, we increase the number of concurrent invocations of the same function (horizontal axes), whereas, for asynchronous invocation, concurrency depends on the system implementation which cannot be forced directly; thereby, we use the request arrival rate to increase the concurrency of invocations. To remove the impact of any randomness, We disabled the auto-scaling and limited the number of worker instances to 6. We also exclude OpenWhisk from this section because the Python runtime in OpenWhisk does not support container-level concurrency.
For the transcoding function (Figures 8a and 8d), at the low concurrency levels ($<$ 80 invocations), Oparaca has average response times higher than Knative, but for the higher concurrency levels, the response time of Knative grows faster than Oparaca due to computing resource limitations. Oparaca doesn’t need to fetch video file metadata, giving it an edge at high concurrency. In the concatenation function (Figures 8b and 8e), however, this phenomenon does not happen. The difference is that text concatenation is IO-intensive and desires high network bandwidth. The overhead of unstructured data access overwhelms the performance gain from eliminating structured data fetching.
For the JSON update function (Figures 8c and 8f), Oparaca can effectively reduce the latency by eliminating the need to fetch from the database. In Figure 8f, because Statefun also shares this invocation scheme and, therefore, offers less completion time than Knative. However, since it relies on local storage to keep the state, while Oparaca uses the memory, Statefun’s completion time is higher than Oparaca’s.
In sum, Oparaca improves performance by eliminating database fetching but adds overhead by accessing unstructured data for secure state abstraction. Depending on the workload, this can either improve or impair object function invocation performance. The overhead may outweigh I/O-intensive workloads, but Oparaca can improve latency by up to 2.27x compared to Knative for workloads without unstructured data.
Takeaway: Object abstraction can be provided with an insignificant latency overhead for objects with only a structured state. The main object overhead occurs as a result of securing unstructured data access.
VI-C Scalability of the Oparaca Platform
To study the scalability, we scale out the Kubernetes workers from 3—12 VMs, each with 16 vCPU cores (in total 48—192 vCPUs). We measured throughput and speedup metrics, focusing on the JSON update function, which does not rely on slow object storage, which becomes the bottleneck of this experiment. We measure the throughput by continually increasing the concurrency until the throughput stops growing (Figure 9b). We assume three VMs as the base speedup=1, and the speedup of other cluster size is calculated with respect to the base value. Moreover, we add two other versions of Oparaca: first is oprc-bypass that uses a standard Kubernetes deployment as its underlying function execution instead of Knative; Second is oprc-bypass-nonpersist that does not persist the object data to the database to measure if Oparaca is not bottleneck by the database write operation.
According to Figure 9a, the speedup of Knative plateaus after reaching 6 VMs. After a thorough investigation, we realized that this plateau is attributed to the database write operation throughput bottleneck. Conversely, Oparaca exhibits the potential for higher speedup enhancement due to its reliance on the distributed in-memory hash table to consolidate data for batch write operations. This approach can boost maximum throughput by up to 3.27x when comparing oprc-bypass with knative.
Figure 9b shows that oprc-bypass yields a higher throughput over the baseline Oparaca. This is because Oparaca sends task data through the Knative internal proxy to offload the task to Knative. While this setup allows for scale-to-zero functionality, bypassing these components leads to even higher throughput. Furthermore, by disabling the database writing operation, which is the bottleneck, oprc-bypass-nonpersist can achieve even higher throughput. Although there isn’t linear scalability due to the limitations of the database write performance, Oparaca significantly improves maximum throughput compared to traditional FaaS systems.
Takeaway: In addition to offering a higher-level abstraction, Oparaca can improve the throughput and response time of its underlying Knative engine via reducing database operations, thereby, mitigating its bottleneck.
VI-D Performance of Localized Locking
In order to analyze the effectiveness of localized locking, we created an alternative version of Oparaca called oprc-cl that is cluster-wide locking. We then conducted an evaluation using a cluster of 12 Invokers and increasing the request arrival rate to measure the overhead of the locking mechanism. To generate requests involving the locking mechanism, we created multiple requests targeting the same object. From Figure 10, the overhead of localized locking remains mostly constant, whereas the overhead of cluster-wide locking rises with higher request rates. However, the cluster-wide version does not display this behavior, as the network communication overhead limits the throughput and hinders high invocation rates.
VI-E Case Study: Development Efficiency Using OaaS
Figure 11: The use case of developing a face expression recognition workflow for an input video.
In this part, we provide a real-world use case of object development using OaaS and its FaaS counterpart and then demonstrate how OaaS makes the development process of cloud-native serverless applications easier and faster. The use case is a video processing application that performs face detection and facial expression recognition. Figure 11 shows the automatic system uploads the video file to the object storage to be processed by the workflow of functions. The workflow includes: Func_1 that splits the video into multiple image segments; Func_2 that detects the face on each sample image frame; and Func_3 performs the facial expression recognition on the detected face image and generates label data in the JSON format. These functions must persist their output object so that the next function in the workflow can consume it.
FaaS implementation. The developer must implement the following steps: (i) Configuring cloud-based object storage and managing access tokens. (ii) Implementing business logic to respond to trigger events. (iii) Manage data within the functions that involve three steps that are allocating the storage addresses, authenticating access to the object storage, and performing fetch and upload operations to the allocated addresses. Upon implementing these functions, the developer has to connect them as a workflow via a function orchestrator service. Finally, the dashboard service invokes the workflow upon receiving a user request and collects the results.
OaaS implementation. The developer defines three classes, namely Video, Image, and Expression, in the form of the three following classes: (a) Video class with frame_extract() functions; and a macro function, df_exp_recog(detect_interval), that includes the whole dataflow of function calls, with the requested sampling period as its input, and an expression_data object as the output. (b) Image class with the face_detect() and exp_recognize() functions. (c) Expression class that does not require any function. The dashboard service calls the wf_exp_recognize(detect_interval) dataflow function directly using the object access interface and receives the Expression object as the output. We note that the developer does not need to be involved in the data locating and authentication steps when developing the class functions because of the abstraction that OaaS provides.
Takeaway: The OaaS paradigm aggregates the state storage and the function workflow in the object abstraction and enables cloud-native dataflow programming. Thus, developers are relieved from the burden of state management, learning the internal mechanics of the functions and pipelining them.
VII Conclusions
In this research, we presented the OaaS paradigm that aimed at simplifying state and workflow management for cloud-native applications. Our prototype, Oparaca, supports both structured and unstructured data types, ensuring consistency through fail-safe state transitions. It also provides secure and low-overhead management for unstructured data using presigned URLs and redirection mechanisms. For structured data, it employs the pure function scheme to transparently manage application data to the developer code and using DHT and consistent hashing to scalably cache the object data and improve the data locality. To make the Oparaca fault-tolerant, we developed the exactly-once and localized locking schemes. To support cloud-native workflow, Oparaca enables declarative dataflow abstraction that hides the concurrency and synchronization concerns from the developer’s perspective. The evaluation results demonstrate that Oparaca streamlines cloud-native programming and is ideal for use cases that require persisting the state or defining a workflow. Oparaca offers scalability with negligible overhead, particularly for compute-intensive tasks. In the future, we plan to develop Oparaca to support application deployment across multiple data centers, streamlining large-scale application development.